Let 5.0 cal of heat is added to a particular substance. As a result, its volume changes from 50 to 100 cm3 while the pressure remains constant at 1.0 atm. (a) By how much did the internal energy of the substance? (b) If the mass of the substance is 2.0x10-3 mol, find the molar heat capacity at constant volume. (c) Find the molar heat capacity at constant pressure


1
Expert's answer
2021-06-02T13:05:30-0400

a)

according to first law of thermodynamics ,

"\Delta U=Q-W"

= 5 x 4.186 - 1.01 x 105 x (100 - 50) x 10-6

"= 15.88 J"


b) we can calculate  molar heat capacity at constant volume as:


"{ΔU\over n}={15.88\over(2*10^{-3}) }= 7940 J/mol"

c) we can calculate molar heat capacity at constant pressure as:



"{Q\over n} = {5*4.186\over (2*10^{-3})} = 10465 J/mol"